Compute the following​ probabilities:

If Y is distributed N (-4​,9​), Pr ​(Y > -2​) ​= ? (Round to four decimal places)
Answer #1

Solution :

Given that ,

Y is distributed N (-4​,9​),

mean = \mu = -4

standard deviation = \sigma = 9

P(Y > -2) = 1 - p( Y < -2)

=1- p P[(Y - \mu ) / \sigma < (-2 - (-4)) / 9 ]

=1- P(z < 0.22)

Using z table,

= 1 - 0.5871

= 0.4129